Can you please help me with this problem? I have to find the equation of the line that passes through the point (2,5) and is perpendicular to the line x-2y=5
The final answer has to be in slope-intercept form.

Answer by mananth(16946)

1 x -2 y = 5
Find the slope of this line

-2 y = -1 x + 5
Divide by -2
y = 1 x + -2.5
Compare this equation with y=mx+b
slope m = 1/2

The slope of a line perpendicular to the above line will be the negative reciprocal
m1*m2=-1
The slope of the required line will be -2

m= -2 ,point ( 2 , 5 )
Find b by plugging the values of m & the point in
y=mx+b
5 = -4.00 + b
b= 9
m= -2
The required equation is y =-2 x + 9
m.ananth@hotmail.ca